A fish looking up through the water sees that the outside world is contained in a circular horizon. If the refractive index of water is 4/3 and th fish is 12cm below the surface of water, find the radius of the circle.

Grade:10

1 Answers

Vikas TU
7 years ago
After makig the figure U would get the final relation as
tanthetha = r/h
where thetha is the angle making with height h (perpendicular and the hypertenuse.)
and this angle is also the TIR angle.
Therfore from,
sinthetha = 1/u = ¾
thanthetha = 3/root(7)
and r = 36/root(7)